# Co-Writing with AI, on Human Terms: Aligning Research with User Demands Across the Writing Process

### v1

@@ Abstract.
As generative AI tools like ChatGPT become integral to everyday writing, critical questions arise about how to preserve writers’ sense of agency and ownership when using these tools. Yet, a systematic understanding of how AI assistance affects different aspects of the writing process–and how this shapes writers’ agency–remains underexplored. To address this gap, we conducted a systematic review of 109 HCI papers using the PRISMA approach. From this literature, we identify four overarching design strategies for AI writing support–structured guidance, guided exploration, active co-writing, and critical feedback–mapped across the four key cognitive processes in writing: planning, translating, reviewing, and monitoring. We complement this analysis with interviews of 15 writers across diverse domains. Our findings reveal that writers’ desired levels of AI intervention vary across the writing process: content-focused writers (e.g., academics) prioritize ownership during planning, while form-focused writers (e.g., creatives) value control over translating and reviewing. Writers’ preferences are also shaped by contextual goals, values, and notions of originality and authorship. By examining when ownership matters, what writers want to own, and how AI interactions shape agency, we surface both alignment and gaps between research and user needs. Our findings offer actionable design guidance for developing human-centered writing tools for co-writing with AI, on human terms.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: PACMHCIpublicationmonth: 8conference: Proceedings of the ACM on Human-Computer Interaction; October 18–22, 2025; Bergen, Norwayjournalyear: 2025doi: 10.1145/XXXXXXX.XXXXXXXccs: Information systems Systematic reviewsccs: Human-centered computing Collaborative and social computingccs: Computing methodologies Artificial intelligence Natural language processingccs: Human-centered computing Interaction design Empirical studies in HCI
@@ 2.1. Theories on Writing Processes
Our analysis employs the Flower and Hayes (1981) model as our primary theoretical lens for several reasons. First, it provides a comprehensive process model that describes writing behaviours. Second, its processes provide an analytical framework sufficient for examining the collaborative writing process between humans and AI systems. While Nystrand’s model analyzes the social relationship between writer and reader, our research focuses on the collaborative interactions during writing, i.e. how humans and AI jointly engage in planning, translating, reviewing, and monitoring processes. The Flower and Hayes model allows us to examine how these cognitive processes are distributed and negotiated between human writers and AI systems during composition. Compared to other writing process theories, the Flower and Hayes model maintains an optimal balance between sophistication and analytical utility for our specific research context.
@@ 2.2. AI-Assisted Writing
Research on AI-assisted writing systems traces back to early implementations focused on creative writing support. Pre-transformers (Vaswani et al., 2017) systems like Creative Help (Roemmele and Gordon, 2015) and Say Anything (Swanson and Gordon, 2012) utilized case-based reasoning and story repositories to generate context-aware sentence suggestions. Clark’s (Clark et al., 2018) work examining user experiences with AI writing prototypes revealed that while participants found AI collaboration satisfying, the resulting text quality did not surpass that of unaided human writers. These early systems laid the groundwork for understanding both the potential and limitations of AI writing assistance.
@@ 2.3. Agency and Ownership in AI-Assisted Writing
Several studies have identified factors that influence users’ sense of agency and ownership in AI writing systems. Kobiella et al (Kobiella et al., 2024) found that participants who viewed AI as an enhancement tool rather than a replacement reported stronger feelings of accomplishment and ownership, while those who perceived their contributions as minimal experienced diminished ownership. Rezwana et al.’s (Rezwana and Maher, 2023) work highlighted that ownership perceptions depend on both contribution levels and leadership in the writing process, suggesting that interaction designs that maximize user agency can enhance ownership. These findings indicate that maintaining user agency and ownership requires careful consideration of interaction design, user control mechanisms, and the balance between AI support and user autonomy.
These investigations have demonstrated the need for continued focus on users’ senses of agency and ownership when writing with AI. However, there are currently no broad reviews of the AI-assisted writing research landscape that have evaluated HCI researchers’ and system designers’ strategies against users’ needs for preserving their agency and ownership throughout their writing process.
@@ 3.1.1. Query Construction
We developed our search query through an iterative process, beginning with a set of seed papers and expert knowledge in the field. We expanded our initial keyword list through multiple refinement cycles. The final search query combined writing-related terms with AI-related terms in order to capture as many potentially-relevant papers as possible: {mdframed}[backgroundcolor=gray!5, linecolor=black]
@@ 3.2. Study 1 Findings
Organizing (Dang et al., 2022; Afrin et al., 2021; Resch and Yankova, 2019; Taiye et al., 2024; Sun et al., 2024; Choe et al., 2024; Shibani et al., 2024; Shaer et al., 2024; Wang et al., 2024) (Kreminski et al., 2020; Schmitt and Buschek, 2021; Wang et al., 2022; Chung et al., 2022; Yuan et al., 2022; Xu et al., 2024; Ghajargar et al., 2022; Kim et al., 2023a; Biermann et al., 2022; Jahanbakhsh et al., 2022; Singh et al., 2023; Wan et al., 2024) (Kim et al., 2023b, a; Ding et al., 2023; Jahanbakhsh et al., 2022) (Wambsganss et al., 2020; Shin et al., 2022; Wang et al., 2022; Zhang et al., 2023b; Lin et al., 2024; Kim et al., 2024b; Jakesch et al., 2023; Poddar et al., 2023) (Goodman et al., 2022; Di Fede et al., 2022; Suh et al., 2024; Reza et al., 2024; Lin et al., 2024)
@@
Reviewing Evaluating (Wambsganss et al., 2022; Shen et al., 2023; Schmidt, 2020; Wambsganss et al., 2021; Ito et al., 2023; Pereira and Barcina, 2019; Taiye et al., 2024; Benharrak et al., 2024; Singh et al., 2024b; Choe et al., 2024; Allen et al., 2018; Li and Wang, 2024; Shibani et al., 2024; Singh et al., 2024a; Rawat et al., 2024; Wang et al., 2024) (Peng et al., 2023; Wang et al., 2022; Hoque et al., 2024; Xu et al., 2024; Yanardag et al., 2021; Gero et al., 2023; Liapis et al., 2023; Booten and Gero, 2021; Wu et al., 2024; Cheng et al., 2024) (Nouri et al., 2023; Robertson et al., 2021; Neyem et al., 2024) (Wambsganss et al., 2022, 2020; Peng et al., 2020; Wang et al., 2022; Kim et al., 2022; Benharrak et al., 2024; Robertson et al., 2021; Cheng et al., 2024) (Cila, 2022)
@@ 3.2.1. Writing Context Characteristics
    Academic (31 Papers). The Academic writing context includes papers that are focused on research, analysis, or educational use. Papers in this context include topics such as assistance with literature review (Choe et al., 2024; Wang et al., 2024), peer review (Neshaei et al., 2024; Sun et al., 2024), academic writing (Taiye et al., 2024; Singh et al., 2024b; Buruk, 2023; Shibani et al., 2024), and essay writing (Dang et al., 2022; Wambsganss et al., 2021; Afrin et al., 2021; Resch and Yankova, 2019; Benharrak et al., 2024). AI-assisted writing systems in this context are often focused on structured skill development for the users. Support for planning processes are typically intended to help users connect and structure their own ideas to accomplish complex tasks (Resch and Yankova, 2019; Taiye et al., 2024). Translating support provides preliminary drafts or helps the user to restructure their ideas in a different form (e.g. point form to prose), but encourages the user to write and integrate ideas on their own (Sun et al., 2024; Singh et al., 2024b; Choi et al., 2024). Reviewing support is delivered in qualitative form, such as suggestions or summaries (Shen et al., 2023; Benharrak et al., 2024; Singh et al., 2024b). Systems do not revise the user’s text directly, instead recommending improvements to prompt the user to revise the work themselves.
@@
    Personal (24 Papers). The Personal writing context concerns self-expression and sharing one’s thoughts, feelings, and experiences. Compared to other contexts, it embraces informality, subjectivity, and authenticity. Writing tasks in this context include non-academic opinion essay writing (Arakawa et al., 2023; Wambsganss et al., 2020; Lee et al., 2022; Zhang et al., 2023b; Dhillon et al., 2024; Li et al., 2024), blog or social media posts (Shin et al., 2022; Cremaschi et al., 2023; Lin et al., 2024; Cai et al., 2024; Janaka et al., 2024; Benharrak et al., 2024; Jakesch et al., 2023; Poddar et al., 2023), personal messages (Kim et al., 2019) and journalling (Kim et al., 2024b). AI support in this context is generally targeted at lay users, emphasizing ease-of-use. Planning and translating support are intermingled due to the frequent use of longer AI outputs that directly ideate and write for the user, though these are generally presented as suggestions in order to preserve the user’s engagement with the text (Cai et al., 2024; Kim et al., 2024b). We also see transformation of user inputs between modalities like speech to text or visuals to text (Zhang et al., 2023b; Lin et al., 2024), or between textual forms like keywords to prose (Kim et al., 2024b). Reviewing support is typically provided through quantitative feedback, with a focus on evaluation rather than direct revision (Wambsganss et al., 2022; Peng et al., 2020).
@@ 4.1.2. Procedure
Each study session lasted between 60 and 90 minutes and was conducted online via recorded video calls by the lead author, allowing us to reach participants across multiple geographic locations. We introduced participants to the study and then asked them to complete a 5-minute pre-survey to provide consent and share demographic information, their writing experience, and AI usage. We informed participants of their right to withdraw from the study at any time and compensated each participant with 20 CAD for their time. The institution’s research ethics board approved the study protocol.
@@ 4.1.3. Data Analysis
The data included transcripts of the interview recordings and responses to pre- and post-interview surveys. To identify factors that influence writers’ sense of ownership in the AI-assisted writing process, we conducted a reflexive thematic analysis (Braun et al., 2023) of transcripts through an inductive-deductive approach. Guided by the cognitive process theory of writing, we used the main writing processes—planning, translation, reviewing, and monitoring—as predefined codes to structure our interpretation, while also inductively identifying new patterns. The pre-survey data provided important context about each participant’s background in writing and prior experience with AI tools. The post-interview survey helped quantify attitudes toward AI across different cognitive processes. Scores for negatively worded items (Q1, Q3, Q5, Q9, and Q13) were reversed (see Figure 4). Given the varied perspectives on ownership across writing elements, our goal was not to aggregate results into a single measure of ownership and agency but rather to examine distinct aspects of the writing process.
@@ 4.2.1. When Ownership Matters
Goal-setting Writer defines key objectives and frameworks; AI used for background structure alignment. W13: “If I outline the goals clearly, ChatGPT can help format, but the primary direction remains mine.” Creative goals set by writer; AI supports structure adjustments. W7: “My voice is in the goals of the story, AI only aids in structure refinement without altering intent.”
@@
Legend: Red - Little Support Yellow - Moderate Support Green - Significant Support
@@ 4.2.3. How AI Interactions Shape Ownership
These examples show how form-oriented writers prioritize their own stylistic and creative sensibilities, even in scenarios with established standards or grammar. They also point to the role that AI interaction design plays in supporting writers with monitoring and decision-making. W8 expands on this idea: I think for me, writing is so much about decision making that’s like what you’re doing at every single stage. And so I think that that’s part of why I feel so attached to the AI being the one that’s suggesting, but not necessarily the one that’s directly editing anything that you’re working on, …so it’s important that those decisions are primarily made by you and not by the AI.”
@@ 5. Alignment Between the Two Studies
Goal-setting Writer defines key objectives and frameworks; AI used for background structure alignment Creative goals set by writer; AI supports structure adjustments Scaffolding of AI system provides pre-defined objectives that must be followed by the user User defines goals, with AI assistance through iterative exploration and selection of ideas AI works collaboratively towards writer-defined goals with some autonomy User maintains control over the text’s goals; AI supports user goals through critical feedback
@@
Legend: Red - Little Support; Yellow - Moderate Support; Green - Significant Support
@@ 5.1. Alignment with Contextual Factors of Ownership
The CSCW community has addressed several of these factors. Shakeri et al. (Shakeri et al., 2021) designed an AI system to enable human-human collaborative writing by offloading narrative tasks to AI. By ceding ownership of narration to AI, while retaining control over their character’s dialogue, users were able to alleviate time constraints and vulnerability caused by a lack of perceived confidence in creative writing. Hauptman et al. (Hauptman et al., 2022) found that professionals’ desire to collaborate with AI was associated with the provision of explainable, actionable feedback and shared social context to build trust, reflecting the preferences of our interview participants. Beyond the text modality, Zhang et al. (Zhang et al., 2023a) created a multi-level human-AI co-creation framework that enables users to customize the level of AI assistance based on their perceived needs and time, though with limited observations of the effects on users. Cao et al. (Cao et al., 2023) investigated the impact of time pressure on human decision-making abilities, and the potential for AI support systems to mitigate these effects. Finally, Tang et al. (Tang et al., 2024) found differential usage patterns of image-generating AI between professional and non-professional users driven by perceived competence and level of trust.

### v2

@@ Abstract.
As generative AI tools like ChatGPT become integral to everyday writing, critical questions arise about how to preserve writers’ sense of agency and ownership when using these tools. Yet, a systematic understanding of how AI assistance affects different aspects of the writing process—and how this shapes writers’ agency—remains underexplored. To address this gap, we conducted a systematic review of 109 HCI papers using the PRISMA approach. From this literature, we identify four overarching design strategies for AI writing support–structured guidance, guided exploration, active co-writing, and critical feedback–mapped across the four key cognitive processes in writing: planning, translating, reviewing, and monitoring. We complement this analysis with interviews of 15 writers across diverse domains. Our findings reveal that writers’ desired levels of AI intervention vary across the writing process: content-focused writers (e.g., academics) prioritize ownership during planning, while form-focused writers (e.g., creatives) value control over translating and reviewing. Writers’ preferences are also shaped by contextual goals, values, and notions of originality and authorship. By examining when ownership matters, what writers want to own, and how AI interactions shape agency, we surface both alignment and gaps between research and user needs. Our findings offer actionable design guidance for developing human-centered writing tools for co-writing with AI, on human terms.
AI-Assisted Writing, Human-Centered AI, Large Language Models
$c$s: General and reference Surveys and overviewsccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Interaction designccs: Computing methodologies Artificial intelligence
@@ 2.1. Theories on Writing Processes
Our analysis employs the Flower and Hayes (1981) (Flower and Hayes, 1981) model as our primary theoretical lens for several reasons. First, it provides a comprehensive process model that describes writing behaviours. Second, its processes provide an analytical framework sufficient for examining the collaborative writing process between humans and AI systems. While Nystrand’s model analyzes the social relationship between writer and reader, our research focuses on the collaborative interactions during writing, i.e. how humans and AI jointly engage in planning, translating, reviewing, and monitoring processes. The Flower and Hayes model allows us to examine how these cognitive processes are distributed and negotiated between human writers and AI systems during composition. Compared to other writing process theories, the Flower and Hayes model maintains an optimal balance between sophistication and analytical utility for our specific research context.
@@ 2.2. AI-Assisted Writing
Research on AI-assisted writing systems traces back to early implementations focused on creative writing support. Pre-transformer (Vaswani et al., 2017) systems like Creative Help (Roemmele and Gordon, 2015) and Say Anything (Swanson and Gordon, 2012) utilized case-based reasoning and story repositories to generate context-aware sentence suggestions. Clark’s (Clark et al., 2018) work examining user experiences with AI writing prototypes revealed that while participants found AI collaboration satisfying, the resulting text quality did not surpass that of unaided human writers. These early systems laid the groundwork for understanding both the potential and limitations of AI writing assistance.
@@ 2.3. Agency and Ownership in AI-Assisted Writing
Several studies have identified factors that influence users’ sense of agency and ownership in AI writing systems. Kobiella et al. (Kobiella et al., 2024) found that participants who viewed AI as an enhancement tool rather than a replacement reported stronger feelings of accomplishment and ownership, while those who perceived their contributions as minimal experienced diminished ownership. Rezwana et al.’s (Rezwana and Maher, 2023) work highlighted that ownership perceptions depend on both contribution levels and leadership in the writing process, suggesting that interaction designs that maximize user agency can enhance ownership. Through a thematic analysis of Human-AI co-creation discussion threads on reddit, Xu et al. (Xu et al., 2024a) surfaced the themes of involvement, sense of infringement, and notion of legal agreements as most relevant in developing a sense of ownership. These findings indicate that maintaining user agency and ownership requires careful consideration of interaction design, user control mechanisms, and the balance between AI support and user autonomy.
These investigations have demonstrated the need for continued focus on users’ senses of agency and ownership when writing with AI. Previous work on ownership in HCI, such as Kuzminykh & Cauchard’s framework (Kuzminykh and Cauchard, 2020), has explored ways to develop a theoretical conceptualization of ownership in a traditional HCI context. However, there are currently no broad reviews of the AI-assisted writing research landscape that have evaluated HCI researchers’ and system designers’ strategies against users’ needs for preserving their agency and ownership throughout their writing process.
@@ 3.1.1. Query Construction
We developed our search query through an iterative process, beginning with a set of seed papers and expert knowledge in the field. We expanded our initial keyword list through multiple refinement cycles. The final search query combined writing-related terms with AI-related terms in order to capture as many potentially-relevant papers as possible:
@@ 3.2. Study 1 Findings
Organizing (Dang et al., 2022; Afrin et al., 2021; Resch and Yankova, 2019; Taiye et al., 2024; Sun et al., 2024; Choe et al., 2024; Shibani et al., 2024; Shaer et al., 2024; Wang et al., 2024) (Kreminski et al., 2020; Schmitt and Buschek, 2021; Wang et al., 2022; Chung et al., 2022; Yuan et al., 2022; Xu et al., 2024b; Ghajargar et al., 2022; Kim et al., 2023a; Biermann et al., 2022; Jahanbakhsh et al., 2022; Singh et al., 2023; Wan et al., 2024) (Kim et al., 2023b, a; Ding et al., 2023; Jahanbakhsh et al., 2022) (Wambsganss et al., 2020; Shin et al., 2022; Wang et al., 2022; Zhang et al., 2023b; Lin et al., 2024; Kim et al., 2024b; Jakesch et al., 2023; Poddar et al., 2023) (Goodman et al., 2022; Di Fede et al., 2022; Suh et al., 2024; Reza et al., 2024; Lin et al., 2024)
@@
Reviewing Evaluating (Wambsganss et al., 2022; Shen et al., 2023; Schmidt, 2020; Wambsganss et al., 2021; Ito et al., 2023; Pereira and Barcina, 2019; Taiye et al., 2024; Benharrak et al., 2024; Singh et al., 2024b; Choe et al., 2024; Allen et al., 2018; Li and Wang, 2024; Shibani et al., 2024; Singh et al., 2024a; Rawat et al., 2024; Wang et al., 2024) (Peng et al., 2023; Wang et al., 2022; Hoque et al., 2024; Xu et al., 2024b; Yanardag et al., 2021; Gero et al., 2023; Liapis et al., 2023; Booten and Gero, 2021; Wu et al., 2024; Cheng et al., 2024) (Nouri et al., 2023; Robertson et al., 2021; Neyem et al., 2024) (Wambsganss et al., 2022, 2020; Peng et al., 2020; Wang et al., 2022; Kim et al., 2022; Benharrak et al., 2024; Robertson et al., 2021; Cheng et al., 2024) (Cila, 2022)
@@ 3.2.1. Writing Context Characteristics
    Academic (31 Papers). The Academic writing context includes papers that are focused on research, analysis, or educational use. Papers in this context include topics such as assistance with literature review (Choe et al., 2024; Wang et al., 2024), peer review (Neshaei et al., 2024; Sun et al., 2024), academic writing (Taiye et al., 2024; Singh et al., 2024b; Buruk, 2023; Shibani et al., 2024), and essay writing (Dang et al., 2022; Wambsganss et al., 2021; Afrin et al., 2021; Resch and Yankova, 2019; Benharrak et al., 2024). AI-assisted writing systems in this context are often focused on structured skill development for the users. Support for planning processes are typically intended to help users connect and structure their own ideas to accomplish complex tasks (Resch and Yankova, 2019; Taiye et al., 2024). Translating support provides preliminary drafts or helps the user to restructure their ideas in a different form (e.g., point form to prose), but encourages the user to write and integrate ideas on their own (Sun et al., 2024; Singh et al., 2024b; Choi et al., 2024). Reviewing support is delivered in qualitative form, such as suggestions or summaries (Shen et al., 2023; Benharrak et al., 2024; Singh et al., 2024b). Systems do not revise the user’s text directly, instead recommending improvements to prompt the user to revise the work themselves.
@@
    Personal (24 Papers). The Personal writing context concerns self-expression and sharing one’s thoughts, feelings, and experiences. Compared to other contexts, it embraces informality, subjectivity, and authenticity. Writing tasks in this context include non-academic opinion essay writing (Arakawa et al., 2023; Wambsganss et al., 2020; Lee et al., 2022; Zhang et al., 2023b; Dhillon et al., 2024; Li et al., 2024), blog or social media posts (Shin et al., 2022; Cremaschi et al., 2023; Lin et al., 2024; Cai et al., 2024; Janaka et al., 2024; Benharrak et al., 2024; Jakesch et al., 2023; Poddar et al., 2023), personal messages (Kim et al., 2019) and journaling (Kim et al., 2024b). AI support in this context is generally targeted at lay users, emphasizing ease-of-use. Planning and translating support are intermingled due to the frequent use of longer AI outputs that directly ideate and write for the user, though these are generally presented as suggestions in order to preserve the user’s engagement with the text (Cai et al., 2024; Kim et al., 2024b). We also see transformation of user inputs between modalities like speech to text or visuals to text (Zhang et al., 2023b; Lin et al., 2024), or between textual forms like keywords to prose (Kim et al., 2024b). Reviewing support is typically provided through quantitative feedback, with a focus on evaluation rather than direct revision (Wambsganss et al., 2022; Peng et al., 2020).
@@ 4.1.2. Procedure
Each study session lasted between 60 and 90 minutes and was conducted online via recorded video calls by the lead author, allowing us to reach participants across multiple geographic locations. We introduced participants to the study and then asked them to complete a 5-minute pre-survey to provide consent and share demographic information, their writing experience, and AI usage. We informed participants of their right to withdraw from the study at any time and compensated each participant with CAN\$20 for their time. The institution’s research ethics board approved the study protocol.
@@ 4.1.3. Data Analysis
The data included transcripts of the interview recordings and responses to pre- and post-interview surveys. To identify factors that influence writers’ sense of ownership in the AI-assisted writing process, we conducted a reflexive thematic analysis (Braun et al., 2023) of transcripts through an inductive-deductive approach. Guided by the cognitive process theory of writing, we used the main writing processes—planning, translation, reviewing, and monitoring—as predefined codes to structure our interpretation, while also inductively identifying new patterns. The pre-survey data provided important context about each participant’s background in writing and prior experience with AI tools. The post-interview survey helped quantify attitudes toward AI across different cognitive processes. Scores for negatively worded items (P1, P3, T1, R1, and M1) were reversed (see Figure 4). Given the varied perspectives on ownership across writing elements, our goal was not to aggregate results into a single measure of ownership and agency but rather to examine distinct aspects of the writing process.
@@ 4.2.1. When Ownership Matters
Goal-setting Writer defines key objectives and frameworks; AI used for background structure alignment. W13: “If I outline the goals clearly, ChatGPT can help format, but the primary direction remains mine.” Creative goals developed by writer; AI provides structure and adjustments. W7: “[AI] does help me understand the shifting goal posts that are moving around at a given moment.”
@@
Legend: Pink - Little Support Yellow - Moderate Support Blue - Significant Support
@@ 4.2.3. How AI Interactions Shape Ownership
These examples show how form-oriented writers prioritize their own stylistic and creative sensibilities, even in scenarios with established standards or grammar. They also point to the role that AI interaction design plays in supporting writers with monitoring and decision-making. W8 expands on this idea: I think for me, writing is so much about decision making that’s like what you’re doing at every single stage. And so I think that that’s part of why I feel so attached to the AI being the one that’s suggesting, but not necessarily the one that’s directly editing anything that you’re working on…so it’s important that those decisions are primarily made by you and not by the AI.”
@@ 5. Alignment Between the Two Studies
Goal-setting Writer defines key objectives and frameworks; AI used for background structure alignment Creative goals developed by writer; AI supports structure adjustments Scaffolding of AI system provides pre-defined objectives that must be followed by the user User defines goals, with AI assistance through iterative exploration and selection of ideas AI works collaboratively towards writer-defined goals with some autonomy User maintains control over the text’s goals; AI supports user goals through critical feedback
@@
Legend: Pink - Little Support; Yellow - Moderate Support; Blue - Significant Support
@@ 5.1. Alignment with Contextual Factors of Ownership
Several previous works within the CSCW community have addressed these factors within social and collaborative contexts. This body of work provides particular insight into how ownership concerns manifest when writing involves multiple stakeholders, shared authorship, or professional collaboration - contexts where questions of agency and ownership become especially complex. Time constraints, a key concern for our participants, have been explored by Shakeri et al. (Shakeri et al., 2021), who found that offloading narrative tasks to AI helped alleviate time pressures in collaborative writing, and by Cao et al. (Cao et al., 2023), who investigated how AI support systems can mitigate the effects of time pressure on decision-making. Trust-building in collaborative environments appears prominently in Hauptman et al. (Hauptman et al., 2022), who found that professionals required explainable, actionable feedback and shared social context to build trust with AI collaborators, reflecting preferences similar to those of our interview participants. Perceived competence differences have been documented by Tang et al. (Tang et al., 2024), who found differential usage patterns between professional and non-professional users of AI tools, and by Shakeri et al., who addressed users’ vulnerability from lack of confidence in creative writing contexts. Task importance considerations appear in Zhang et al.’s (Zhang et al., 2023a) multi-level framework, which enables users to choose between different levels of AI creative involvement based on their individualized needs with a design that could accommodate varying levels of control desired for different stakes of writing tasks. While individual studies have addressed one or two of these factors, our analysis reveals how all four factors shape writers’ ownership concerns in AI-assisted writing. This framework provides a foundation for future research to consider the full range of contextual influences on AI writing collaboration, rather than examining these factors in isolation.